# Exploring 3D Activity Reasoning and Planning: From Implicit Human Intentions to Route-Aware Planning

### v2

@@ 4.4 Dynamic Graph Modulation
where node $N_{ij}$ denotes the $j$-th neighbor of $N_{i}$, and edge $E_{ij}$ represents the spatial relation between $N_{ij}$ and $N_{i}$. Here, the $K$ neighbors of node $N_{i}$ are selected using the K-Nearest Neighbors (KNN) algorithm. This mechanism can effectively coordinate with the proposed Progressive Plan Generation approach in the iterative process, as illustrated in Algorithm 1.
@@ 4.5 Training Objective
5 Experiment
5.1 Experimental Settings
@@
We adopt BLEU [37], ROUGE [32], METEOR [3], and CIDEr [49] to evaluate the quality of generated plans for the reasoned activities. In addition, we adopt BLEU-4, CIDEr, and METEOR to measure the performance of activity reasoning. Since our task centers on human-robot interaction and textual response generation, particularly for activity reasoning and planning, we adopt language-based metrics to evaluate the overall quality of the responses.
@@
We conduct experiments on one NVIDIA A40 GPU (48 GB memory), and train our framework for $11k$ iterations with a batch size of $2$ and a total training time of around 9 hours. We employ the AdamW [33] optimizer with an initial learning rate of $2e^{-5}$, a weight decay of $0.02$, and a cosine scheduler with $1.1k$ warm-up iterations. We adopt LLaMA-3-8B-Instruct [1] as our multimodal LLM backbone and apply LoRA [19] with a rank of 16 for efficient fine-tuning. The weight $w_{l}$ used in the proposed Dynamic Graph Modulation is set as $2.0$.
Table 2: Benchmarking on the ReasonPlan3D validation set for the 3D reasoning and planning task with evaluation metrics BLEU, CIDEr, METEOR, and ROUGE. Best in bold, second underlined.
Method Venue BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr METEOR ROUGE
3D-LLM [18] NeurIPS 23 34.47 25.73 21.59 16.52 10.02 18.61 40.29
LL3DA [5] CVPR 24 38.21 28.92 22.74 19.86 10.93 21.55 40.72
Chat-Scene [20] NeurIPS 24 39.47 31.04 25.80 21.15 11.67 21.94 42.28
3DGraphLLM [56] arXiv 24 40.29 31.37 25.37 20.98 11.82 22.51 43.60
SHARP (Ours) - 52.22 39.30 30.58 24.59 25.51 27.51 43.76
5.2 Benchmarking with Existing Methods
@@
In addition, Table 3 presents the activity reasoning performance of SHARP and several state-of-the-art methods over the validation set of ReasonPlan3D. Our approach achieves superior performance across all metrics. Specifically, 3D-LLM [18], LL3DA [5], and Chat-Scene [20] do not employ a scene graph to explicitly guide the model’s awareness of critical objects, leading to suboptimal performance when dealing with the activity reasoning task where comprehending object relations plays an essential role. 3DGraphLLM [56] employs only scene graphs without scene-level embeddings of the global picture, leading to inferior 3D scene understanding and weaker performance for activity reasoning.
@@
      Method       BLEU-4       CIDEr       METEOR
      3D-LLM [18]       38.02       285.46       30.13
      LL3DA [5]       42.59       302.47       33.20
      Chat-Scene [20]       45.26       313.90       34.42
      3DGraphLLM [56]       47.65       329.59       36.87
      SHARP (Ours)       50.28       359.53       38.35
Qualitative Benchmarking.
Figure 5 provides qualitative results on the validation set of ReasonPlan3D. Each example presents an implicit human instruction, an input point cloud, the corresponding 3D scene segmentation, and the generated textual answers by 3DGraphLLM [56] and the proposed SHARP. We can observe that SHARP can reason activities from the implicit input instructions accurately. Moreover, SHARP performs well in generating high-quality route plans as well. In both examples, SHARP surpasses 3DGraphLLM by generating more detailed steps, demonstrating its superior capability in complex planning tasks. Moreover, 3DGraphLLM introduces objects that do not exist in the scene, such as “couch”, “cushions”, and “coffee table” in the first example, while SHARP exhibits a more precise understanding of the 3D environment, generating plans that accurately reflect the objects present in the scene.
@@
Table 5: Ablation studies on the weight $w_{l}$ as used in the proposed Dynamic Graph Modulation. The best results are in bold.
Index $w_{l}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr METEOR ROUGE
1 1.0 51.46 38.49 29.35 23.82 23.80 27.09 43.39
2 1.5 51.87 38.84 30.11 24.18 24.75 27.36 43.52
3 2.0 52.22 39.30 30.58 24.59 25.51 27.51 43.76
4 3.0 51.96 39.01 30.16 24.20 25.09 27.38 43.59
5.3 Ablation Study
We conduct ablation studies over the validation set of ReasonPlan3D to evaluate the effectiveness of our designs, including Progressive Plan Generation and Dynamic Graph Modulation as presented in the SHARP framework. In addition, we examine the effect of the weight $w_{l}$ as used in the Dynamic Graph Modulation.
@@ 5.3 Ablation Study
Effectiveness of $w_{l}$.
As defined in Equation 1, the weight $w_{l}$ in the Dynamic Graph Modulation is introduced to highlight critical objects and their spatial relations relevant to the reasoned activity. We examine the effectiveness of $w_{l}$ in Table 5. It can be observed that critical objects and their spatial relations are not well captured when $w_{l}$ is set as $1.0$, leading to degraded performance. Increasing $w_{l}$ to 1.5 consistently improves performance, suggesting that placing greater emphasis on critical objects and their spatial relations can boost the model’s ability. Further increasing $w_{l}$ to 2.0 achieves the best performance, confirming that guiding the model toward these crucial cues enhances the model’s planning ability. However, increasing $w_{l}$ to 3.0 overemphasizes these objects and their spatial relations, leading to an imbalance among objects in the 3D scene that degrades the performance.
6 Conclusion
This paper presents a novel 3D Activity Reasoning and Planning task that reasons activities from implicit human intentions to satisfy their requirements, decomposes tasks into sequential steps, and plans inter-step routes in complex 3D environments. To this end, we introduce ReasonPlan3D, a large-scale benchmark containing diverse implicit user instructions, step-by-step task planning, inter-step route planning, and fine-grained segmentation annotations. On top of this, we propose a novel framework with a progressive plan generation mechanism to ensure that the generated steps remain contextually consistent. Moreover, a scene graph for capturing object relations is constructed, which is dynamically updated to emphasize critical objects and spatial relations for improving route planning accuracy. Extensive experiments confirm the effectiveness of our approach. Future work will investigate more challenging 3D scenarios, further expanding the potential of 3D scene understanding and human-robot interaction in real-world applications.

### v3

@@ Introduction
where node $N_{ij}$ denotes the $j$-th neighbor of $N_{i}$, and edge $E_{ij}$ represents the spatial relation between $N_{ij}$ and $N_{i}$. Here, $K$ neighbors of node $N_{i}$ are selected by the K-Nearest Neighbors (KNN) algorithm. This mechanism can effectively coordinate with the proposed Progressive Plan Generation approach in the iterative process, as shown in Figure 3.
Method Venue BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr METEOR ROUGE
3D-LLM (Hong et al. 2023) NeurIPS 23 34.47 25.73 21.59 16.52 10.02 18.61 40.29
LL3DA (Chen et al. 2024a) CVPR 24 38.21 28.92 22.74 19.86 10.93 21.55 40.72
Chat-Scene (Huang et al. 2024a) NeurIPS 24 39.47 31.04 25.80 21.15 11.67 21.94 42.28
3DGraphLLM (Zemskova and Yudin 2024) arXiv 24 40.29 31.37 25.37 20.98 11.82 22.51 43.60
SHARP (Ours) - 52.22 39.30 30.58 24.59 25.51 27.51 43.76
Table 2: Benchmarking on the ReasonPlan3D validation set for the 3D Reasoning-Driven Planning task with evaluation metrics BLEU, CIDEr, METEOR, and ROUGE. Best in bold, second underlined.
    Method     BLEU-4     CIDEr     METEOR
    3D-LLM (Hong et al. 2023)     38.02     285.46     30.13
    LL3DA (Chen et al. 2024a)     42.59     302.47     33.20
    Chat-Scene (Huang et al. 2024a)     45.26     313.90     34.42
    3DGraphLLM (Zemskova and Yudin 2024)     47.65     329.59     36.87
    SHARP (Ours)     50.28     359.53     38.35
@@
Experiment
Experimental Settings
@@
We adopt BLEU (Papineni et al. 2002), ROUGE (Lin 2004), METEOR (Banerjee and Lavie 2005), and CIDEr (Vedantam, Lawrence Zitnick, and Parikh 2015) to evaluate the quality of generated plans for the reasoned activities. In addition, we adopt BLEU-4, CIDEr, and METEOR to measure the performance of activity reasoning. Since our task centers on human-robot interaction and textual response generation, particularly for activity reasoning and planning, we adopt language-based metrics to evaluate the overall quality of the responses.
@@
[…]
Benchmarking with Existing Methods
@@
In addition, Table 3 presents the activity reasoning performance of SHARP and several state-of-the-art methods over the validation set of ReasonPlan3D. Our approach achieves superior performance across all metrics. Specifically, 3D-LLM (Hong et al. 2023), LL3DA (Chen et al. 2024a), and Chat-Scene (Huang et al. 2024a) do not employ a scene graph to explicitly guide the model’s awareness of critical objects, leading to suboptimal performance when dealing with the activity reasoning task, where comprehending object relations plays an essential role. 3DGraphLLM (Zemskova and Yudin 2024) employs only scene graphs without scene-level embeddings, leading to inferior 3D scene understanding and weaker performance for activity reasoning.
@@
Qualitative Benchmarking.
Figure 4 provides qualitative results on the validation set of ReasonPlan3D. Each example presents an implicit human instruction, an input point cloud, the corresponding 3D scene segmentation, and the generated textual answers by 3DGraphLLM (Zemskova and Yudin 2024) and the proposed SHARP. We can observe that SHARP can reason activities from the implicit input instructions accurately while generating high-quality route plans. In both examples, SHARP surpasses 3DGraphLLM by generating more detailed steps, demonstrating its superior capability in complex planning tasks. Moreover, 3DGraphLLM introduces objects that do not exist in the scene, such as “couch”, “cushions”, and “coffee table” in the first example, while SHARP exhibits a more precise understanding of the 3D environment, generating plans that accurately reflect the objects present in the scene.
Ablation Study
We conduct ablation studies over the validation set of ReasonPlan3D to evaluate the effectiveness of our designs, including Progressive Plan Generation and Dynamic Graph Modulation in the SHARP framework.
@@
Conclusion
This paper presents a novel 3D Reasoning-Driven Planning task that reasons activities from implicit human intentions to satisfy their requirements, decomposes tasks into sequential steps, and plans inter-step routes in complex 3D environments. To this end, we introduce ReasonPlan3D, a large-scale benchmark containing diverse implicit user instructions, step-by-step task planning, inter-step route planning, and fine-grained segmentation annotations. On top of this, we propose a novel framework with a progressive plan generation mechanism to ensure that the generated steps remain contextually consistent. Moreover, a scene graph for capturing object relations is constructed, which is dynamically updated to emphasize critical objects and spatial relations for improving route planning accuracy. Extensive experiments confirm the effectiveness of our approach. Future work will investigate more challenging 3D scenarios, further expanding the potential of 3D scene understanding and human-robot interaction in real-world applications.